If the length of a rectangle is 3 feet longer than the width and
the diagonal is 15 feet, then what are the length and width?

Answer by Greenfinch(383)

width is x, length is x + 3, diagonal is 15
using Pythagoras x^2 +(x + 3)^2 = 225
2x^2 + 6x + 9 = 225
2x^2 + 6x -216 = 0
x^2 + 3x - 108 = 0
(x + 12 )(x - 9)= 0
x = -12, or +9
width is 9, length is 12, diagonal is 15 (a 3,4,5 triangle)